OpenStudy (anonymous):

Write the equation of the line which passes through (2, –3) and is perpendicular to y = 4x + 7 in standard form. (2 point)

OpenStudy (anonymous):

perperndicular lines have reciprocal and opposite side of the given equation so; for (2;-3) m=-1/4 then y=-1/4(x)+c -3=-1/4(2)+c c=-5/2 then y=-1/4(x)-5/2 :this is the equation

OpenStudy (anonymous):

in standard form it will be 1/4(x)+y = -5/2